OpenStudy (anonymous):

The price-demand equation for heads of lettuce is 40p-x=25 . Where P is the price of a head of lettuce and X is the weekly demand (in thousands) for lettuce. How can I write an expression for revenue as a function of the weekly demand for lettuce.

OpenStudy (gw2011):

Since x is the weekly demand and this is what you are looking for, then you can solve the equation for x: 40p-x=25 40p-25=x

OpenStudy (anonymous):

how

OpenStudy (gw2011):

If you know what the actual price (p) is, then you substitute that into x=40p-25 and thiswill give you the weekly demand.

OpenStudy (anonymous):

i tried it but its incorrect.

OpenStudy (gw2011):

Revenue=(Lettuce Sold)(Price) which is (P)(X). I guess I'm not understanding the equation 40p-x=25. According to this equation it seems to me that if you are looking for x which is the demand for lettuce (lettuce sold), then the equation is: x=40p-25. If you want the price, then the equation is: p=(x+25)/40. I don't know if this will help you.

OpenStudy (anonymous):

i just need to write it as a revenue function. i dont nee to solve for anything...i think.